# Hyperuniformity and optimal transport of point processes

### v3

@@ 3.2 Proof of Theorems 3 and 4
For $k=mn^{-1/d}$
$\mathbf{E}\left[\frac{1}{N}S_{{n}}(k)\right]=$ $\mathbf{E}\frac{1}{N^{2}}\left|\sum_{x\in\mu_{n}}e^{-ik\cdot x}\right|^{2}\leqslant 4\mathbf{E}\left(\frac{1}{n^{2}}\left|\sum_{x{\in\mu_{n}}}e^{-ik\cdot x}\right|^{2}\right)+\mathbf{E}\left(\frac{1}{N^{2}}N^{2}\mathbf{1}_{\{N<n/2\}}\right).$ (37)
@@
$\mathbf{P}(N<n/2)\leqslant\mathbf{P}(|N-\mathbf{E}(N)|>n/2)\leqslant\frac{4\text{\rm Var}(N)}{n^{2}}$
@@
${\int_{\Lambda_{n}}e^{ik\cdot x}\mathrm{d}x}=0,$
@@
$=$ $1+\int e^{-ik\cdot x}\frac{1}{|\Lambda_{n}|}\int 1_{\Lambda_{n}}(y)\mathrm{d}y\beta(\mathrm{d}x)$
$=$ $1+{n^{-1}}\,\int e^{-ik\cdot x}\,\int{1_{\Lambda_{n}(y)}}1_{\Lambda_{n}}(y+x)\mathrm{d}y\beta(\mathrm{d}x)+n^{-1}\int e^{-ik\cdot x}1_{\Lambda_{n}}(y)1_{\Lambda_{n}^{c}}(y+x)\mathrm{d}y\beta(\mathrm{d}x)$
$=$ $1+{n^{-1}}\,\int e^{-ik\cdot(x+y)}1_{\Lambda_{n}}(x+y)e^{ik\cdot y}1_{\Lambda_{n}}(y)\mathrm{d}y\beta(\mathrm{d}x)+n^{-1}\int e^{-ik\cdot x}\gamma_{n}(x)\beta(\mathrm{d}x)$
$=$ $1+{n^{-1}}\,\mathbf{E}(\sum_{x\neq y\in\mu_{n}}1_{x\neq y}e^{ik\cdot(x-y)}\mu(\mathrm{d}x)\mu(\mathrm{d}y))+\underbrace{\left|{\int_{\Lambda_{n}}e^{ik\cdot x}\mathrm{d}x}\right|^{2}}_{=0}+O\left(\varepsilon(n^{-1/d})\right)$
$=$ $\frac{1}{n}\mathbf{E}\left|\sum_{x{\in\mu_{n}}}e^{-ik\cdot x}\right|^{2}+O\left(\varepsilon(n^{-1/d})\right)$
using (3) and (36). In the fourth equality above, we have used that $\gamma_{n}$ and $\beta$ are even functions; the former by definition and the latter due to stationarity of $\mu$. Plugging back into (37) yields (29).
@@
$\mathbf{E}\big{(}\sum_{x\neq y\in\mu_{n}}1_{x\neq y}e^{ik\cdot(x-y)}\mu(\mathrm{d}x)\mu(\mathrm{d}y)\big{)}=$ $\int e^{-ik\cdot(x+y)}1_{\Lambda_{n}}(x+y)e^{ik\cdot y}1_{\Lambda_{n}(y)}\mathrm{d}y\beta(\mathrm{d}x)$
$\leqslant$ $|\Lambda_{n}|\int 1_{\Lambda_{n}+\Lambda_{n}}(x)|\beta|(\mathrm{d}x).$
Hence in the non-integrable case, with Lemma 8
$\mathbf{E}(N^{-1}S_{n}(k))\leqslant\frac{1}{n^{2}}n|\beta|(\Lambda_{n}+\Lambda_{n})+\frac{\text{Var}(N)}{n^{2}}\leqslant c\,\frac{\mathsf{b}_{n}}{n}$
@@
$=$ $|\Lambda_{n}|+\int_{\Lambda_{n}^{2}}\mathbf{1}_{\{x\in\Lambda_{n},x+z\in\Lambda_{n}\}}\mathrm{d}x\beta(\mathrm{d}z)$
In the non-integrable case, the bound is
@@
$=$ $n(1+\beta(\mathbb{R}^{d}))-\int\gamma_{n}(z)\beta(\mathrm{d}z),$ (38)
@@
$\text{Var}(N)\geqslant n(1+\beta(\mathbb{R}^{d})+\varepsilon(n^{-1/d})).$
@@ 3.3 Local to global allocation
The following proposition formalizes and generalizes one of the steps in the proof of [30, Theorem 1(iii)] and is very important to deduce properties of global matchings via properties of matchings in finite windows. Though it is possible to deduce the upcoming Proposition 11 from the more general result in [22, Proposition 2.11], we give a slightly different and self-contained proof.
Recall that for a measure $\mu$, $\tilde{\mu}_{n}=\frac{n}{N}\mu 1_{\Lambda_{n}}$ (with $\tilde{\mu}_{n}=n\delta_{0}$ for $N=0$) and ~$\tilde{\mathcal{L}}_{n}:=\frac{1}{(2\pi)^{d}}\mathcal{L}_{n}$ are the renormalised samples with mass $n$ on $\Lambda_{n}.$
@@
Let $\mu,\nu$ be independent stationary random measures with integrable RPCM and $w:[0,\infty)\to[0,\infty)$ be a continuous increasing weight function. Then the following hold.
  1. 1.
    Assume $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$. Then, there exists a translation invariant coupling $M$ of $\mu$ and $\nu$ whose typical distance $X$ satisfies $\mathbf{E}w(X)<\infty$.
  2. 2.
    Assume that $w(x)\leqslant Cx^{d}$ and $w(x+y)\leqslant C(w(x)+w(y))$ for all $x,y\in[0,\infty)$ and some $C\geqslant 1$. If $\mu,\nu$ are unit intensity simple independent point processes such that $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$ and similarly for $\tilde{\nu}_{n}$, then there is a coupling $M$ of $\mu$ and $\nu$ which is a.s. represented by a one-to-one mapping and whose typical distance $X$ satisfies $\mathbf{E}w(X)<\infty$.
The second point is an infinite version of the following finite sample version, a consequence of the Birkhoff-von Neumann theorem ([50, Exercise 40]).
Proposition 12.
Let $n\geqslant 1$, $\mu=\sum_{i=1}^{{n}}\delta_{x_{i}},\nu=\sum_{i=1}^{n}\delta_{y_{i}}$, where the $x_{i},y_{i}$ are distinct points of $\mathbb{R}^{d}$, and let $w$ a non-decreasing cost function. Then the optimal coupling $M$ is realised by a transport map.
[…]
Lemma 13.
  1. 1.
    If a sequence of random stationary measures $(M_{n})_{n}$ satisfies
    $\sup_{n}\mathbf{E}M_{n}(A)<\infty$
@@
  2. 2.
    Let $\mu_{n},{\nu_{n}}$ be finite (deterministic) atomic measures for $n\geqslant 1$, with the same number of atoms each with mass $a_{n}\to 1$ and also converging resp. vaguely to simple point processes $\mu,{\nu}$. Let $M_{n}$ be a coupling of $\mu_{n},{\nu}_{n}$ that is induced by a transport map $T_{n}:\textrm{supp}(\mu_{n})\to\textrm{supp}({\nu}_{n})$, where supp denotes support of a measure. If $M_{n}\to M$ vaguely, then $M$ is a matching induced by a transport map $T:\mu\to{\nu}$.
Proof.
The first item is a consequence of [36, Theorem 16.15]. So we prove the second item alone here. First, for test functions $f,g\in C_{c}(\mathbb{R}^{d})$, $M(f,1)=\lim_{n}M_{n}(f,1)=\lim_{n}\mu_{n}(f)=\mu(f)$ and similarly $M(1,g)={\nu}(g)$, hence $M$ is indeed a coupling between $\mu$ and ${\nu}$, whose support is contained in $\mu\times{\nu}.$ Since $\mu$ (resp. $\nu$) is assumed to have distinct atoms, for $x\in{\rm supp}(\mu)$, there is $A_{x}$ (resp. $B_{y}$) open containing $x\in{\rm supp}(\mu)$ (resp. $y\in{\rm supp}(\nu)$) such that $A_{x}\cap A_{x^{\prime}}=\emptyset$ (resp. $B_{y}\cap B_{y^{\prime}}=\emptyset$) for $x\neq x^{\prime}$ (resp. $y\neq y^{\prime}$). For $x\in\mu,y\in{\nu}$, vague convergence yields for $n$ large enough, $\mu_{n}(A_{x})={\nu}_{n}(B_{y})=a_{n}$, hence for such large $n$,
$M_{n}(A_{x},B_{y})=a_{n}\mathbf{1}[y=T_{n}(x)],$
and so $M(A_{x},B_{y})=\lim_{n}M_{n}(A_{x},B_{y})\in\{0,1\}$. We finally have for $x\in\mu$
$1=\mu(A_{x})=M(A_{x},\mathbb{R}^{d})=\sum_{y\in{\nu}}\underbrace{M(A_{x},B_{y})}_{\in\{0,1\}}.$
Hence for each $x\in\mu$ there is a unique $y\in{\nu}$ such that $M(A_{x},B_{y})=1$, and we set $T(x):=y$ which indeed represents $M$ as a transport map.
@@
First step: Existence of a translation-invariant infinite coupling $M$:
By assumption, there is a transport plan $m_{n}$ from $\tilde{\mu}_{n}$ to $\tilde{\nu}_{n}$ such that
@@
$m_{n}(\cdot,\mathbb{R}^{d})=\tilde{\nu}_{n}$, $m_{n}(\mathbb{R}^{d},\cdot)=\tilde{\mu}_{n}$ where $d_{n}$ is the toroidal distance on $\Lambda_{n}$. Suppose we tile $\mathbb{R}^{d}$ with randomly shifted copies of $\Lambda_{n}$ and place $\tilde{\mu}_{n},\tilde{\nu}_{n}$ in each of them and translate $m_{n}$ in each copy. We call this transport map $M_{n}$. Formally,
$M_{n}(A\times B)=\sum_{z\in\mathbb{Z}^{d}}m_{n}\big{(}(A-U_{n}+{2\pi}n^{1/d}z)\times(B-U_{n}+{2\pi}n^{1/d}z)\big{)},$ (40)
for $A,B\subset\mathbb{R}^{d}$ and where $U_{n}$ is uniform in $\Lambda_{n}$ and independent of $\mu$ and $\nu$. Observe that $M_{n}$’s are translation invariant, i.e.
$M_{n}(A+x,B+x)\stackrel{{\scriptstyle(d)}}{{=}}M_{n}(A,B)$ (41)
for all $A,B,x$. Note that $M_{n}(A,B)=0$ if $A,B$ are contained in different shifted copies.
Note that $\mathbf{E}M_{n}(A,B)\leqslant\mathbf{E}M_{n}(A,\mathbb{R}^{d})$ and $\mathbf{E}M_{n}(A,\mathbb{R}^{d})$ is translation invariant in $A$. Thus $\mathbf{E}M_{n}(A,\mathbb{R}^{d})=c_{n}|A|$ for some $c_{n}\in[0,\infty]$. For $A=\Lambda_{n}$, we have that $\mathbf{E}M_{n}(A,\mathbb{R}^{d})=\mathbf{E}\tilde{\mu}_{n}(\Lambda_{n})=n$ and so we deduce that $\mathbf{E}M_{n}(A,\mathbb{R}^{d})=|A|$. Similarly, we can deduce that $\mathbf{E}M_{n}(\mathbb{R}^{d},B)=|B|$. Hence, we have that $\mathbf{E}M_{n}(A,B)\leqslant\min(|A|,|B|)$.
In particular, $\mathbf{E}M_{n}(K)<\infty$ for bounded $K\subset\mathbb{R}^{d}\times\mathbb{R}^{d}$. Thus, the sequence $M_{n}$ is tight in vague topology by Lemma 13(1) and has a subsequential distributional limit under vague topology, say $M_{n^{\prime}}\to M$. $M$ is translation invariant as (41) passes to the limit for regular $A,B.$
@@
Denote $\tilde{d}_{n}(x,y)$ to be the distance which is 00 if $x,y$ are in different randomly shifted copies of $\Lambda_{n}$, and the toric distance of the copy otherwise. So $\mathbf{E}[\int\int w(\tilde{d}_{n}(x,y))\mathbf{1}[x\in A]M_{n}(\mathrm{d}(x,y))]$ is translation invariant in $A$ and hence its intensity is either infinity or of the form $\textrm{const}\cdot\mathcal{L}$, and for $A=\Lambda_{n}$, by (39)
@@
Thus we obtain that, for all $A$
@@
Now let $A,B$ be two bounded regular sets i.e. with negligible boundary. Let $n_{0}$ be such that $\sup_{x\in A,y\in B}\|x-y\|<n^{1/d}/2$ for all $n\geqslant n_{0}$. Hence for $x\in A,y\in B$, either $x,y$ are in different shifted copies of $\Lambda_{n}$ and $M_{n}(\mathrm{d}x,\mathrm{d}y)=0$, or they are in the same copy of $\Lambda_{n}$, but at distance $<\frac{n^{1/d}}{2}$, hence $\tilde{d}_{n}(x,y)=\|x-y\|.$ In any case,
@@
$\mathbf{E}\int\int w(\|x-y\|)\mathbf{1}[x\in A,y\in B]M_{n}(\mathrm{d}(x,y))$ $=\mathbf{E}\int\int w(\tilde{d}_{n}(x,y))\mathbf{1}[x\in A,y\in B]M_{n}(\mathrm{d}(x,y))$
$\leqslant c\mathcal{L}(A).$
Since $A,B$ are bounded and have a negligible boundary, and $M$ is invariant under translations of any of its two arguments and has finite intensity, $A\times B$ is a.s. regular for $M$. So, the vague convergence $M_{n}\to M$ easily yields that a.s. $\tilde{w}:(x,y)\mapsto w({\|x-y\|})\mathbf{1}[x\in A,y\in B]$ satisfies
$\int\int\tilde{w}(x,y)\,M_{n}(\mathrm{d}(x,y))\to\int\int\tilde{w}(x,y)\,M(\mathrm{d}(x,y)),$
and hence
$\mathbf{E}\int\int w(\|x-y\|)\mathbf{1}[x\in A,y\in B]M(\mathrm{d}(x,y))\leqslant c\mathcal{L}(A).$
Now letting $B$ increase to $\mathbb{R}^{d}$ and using monotone convergence theorem, shows that the above inequality holds for $B=\mathbb{R}^{d}$ as needed to verify finite typical cost.
@@
We will now show that $M$ is transport plan between independent copies of $\mu$ and $\nu$ defined on a suitable probability space.

### v4

@@ 3.2 Proof of Theorems 3 and 4
Recall that $\pi_{d}:=(2\pi)^{d}$, the volume of the box $\Lambda_{1}$. For $k=mn^{-1/d}$
$\mathbf{E}\left[\frac{1}{N}S_{{n}}(k)\right]=$ $\mathbf{E}\frac{1}{N^{2}}\left|\sum_{x\in\mu_{n}}e^{-ik\cdot x}\right|^{2}\leqslant\frac{4}{\pi_{d}^{2}}\mathbf{E}\left(\frac{1}{n^{2}}\left|\sum_{x{\in\mu_{n}}}e^{-ik\cdot x}\right|^{2}\right)+\mathbf{E}\left(\frac{1}{N^{2}}N^{2}\mathbf{1}_{\{N<\pi_{d}n/2\}}\right).$ (41)
@@
$\mathbf{P}(N<\pi_{d}n/2)\leqslant\mathbf{P}(|N-\mathbf{E}(N)|\geq\pi_{d}n/2)\leqslant\frac{4\text{\rm Var}(N)}{\pi_{d}^{2}n^{2}}$
@@
${\int_{\Lambda_{n}}e^{-ik\cdot x}\mathrm{d}x}=0,$
@@
$=$ $1+\int e^{-ik\cdot x}\frac{1}{|\Lambda_{n}|}\int\mathbf{1}_{\Lambda_{n}}(y)\mathrm{d}y\beta(\mathrm{d}x)$
$=$ $1+\frac{1}{\pi_{d}n}\,\int e^{-ik\cdot x}\,\int{\mathbf{1}_{\Lambda_{n}}(y)}\mathbf{1}_{\Lambda_{n}}(y+x)\mathrm{d}y\beta(\mathrm{d}x)+\frac{1}{\pi_{d}n}\int\int e^{-ik\cdot x}\mathbf{1}_{\Lambda_{n}}(y)\mathbf{1}_{(\Lambda_{n}-x)^{c}}(y)\mathrm{d}y\beta(\mathrm{d}x)$
$=$ $1+\frac{1}{\pi_{d}n}\,\int\int e^{-ik\cdot(x+y)}\mathbf{1}_{\Lambda_{n}}(x+y)e^{-ik\cdot y}\mathbf{1}_{\Lambda_{n}}(y)\mathrm{d}y\beta(\mathrm{d}x)+\frac{1}{\pi_{d}n}\int e^{-ik\cdot x}\gamma_{n}(-x)\beta(\mathrm{d}x)$
$=$ $1+\frac{1}{\pi_{d}n}\,\mathbf{E}\int_{\Lambda_{n}^{2}}\mathbf{1}[x\neq y]e^{-ik\cdot(x-y)}\mu(\mathrm{d}x)\mu(\mathrm{d}y)+\underbrace{\left|{\int_{\Lambda_{n}}e^{-ik\cdot x}\mathrm{d}x}\right|^{2}}_{=0}+O\left(\varepsilon(n^{-1/d})\right)$
$=$ $\frac{1}{\pi_{d}n}\mathbf{E}\sum_{x{\in\mu_{n}}}e^{-ik\cdot(x-x)}+\frac{1}{\pi_{d}n}\mathbf{E}\sum_{x\neq y{\in\mu_{n}}}e^{-ik\cdot(x-y)}+O\left(\varepsilon(n^{-1/d})\right)$
$=$ $\frac{1}{\pi_{d}n}\mathbf{E}\left|\sum_{x{\in\mu_{n}}}e^{-ik\cdot x}\right|^{2}+O\left(\varepsilon(n^{-1/d})\right)$
where we have used (40) in the fifth equality and (5) in the penultimate equality. In the fifth equality above, we have used that $\gamma_{n}$ and $\beta$ are even functions; the former by definition and the latter due to stationarity of $\mu$. Plugging back into (41) yields (32).
@@
$\mathbf{E}\int_{\Lambda_{n}^{2}}\mathbf{1}[x\neq y]e^{-ik\cdot(x-y)}\mu(\mathrm{d}x)\mu(\mathrm{d}y)=$ $\int e^{-ik\cdot(x+y)}\mathbf{1}_{\Lambda_{n}}(y)e^{-ik\cdot y}\mathbf{1}_{\Lambda_{n}}(y+x)\mathrm{d}y\beta(\mathrm{d}x)$
$\leqslant$ $|\Lambda_{n}|\int\mathbf{1}_{\Lambda_{n}+\Lambda_{n}}(x)|\beta|(\mathrm{d}x).$ (42)
Hence in the non-integrable case, recalling $b_{n}$ from (11) and together with Lemma 8
$\mathbf{E}(N^{-1}S_{n}(k))\leqslant\frac{\pi_{d}n}{\pi_{d}n^{2}}\,{|\beta|}(\Lambda_{n}+\Lambda_{n})+\frac{\text{Var}(N)}{\pi_{d}^{2}n^{2}}\leqslant c\,\frac{\mathsf{b}_{n}}{n}$
@@
$=$ $|\Lambda_{n}|+\int\mathbf{1}[x\in\Lambda_{n},x+z\in\Lambda_{n}]\mathrm{d}x\beta(\mathrm{d}z).$
In the non-integrable case, arguing as in (42), we obtain
@@
$=$ $\pi_{d}n(1+\beta(\mathbb{R}^{d}))-\int\gamma_{n}(z)\beta(\mathrm{d}z),$ (43)
@@
$\text{Var}(N)\geqslant\pi_{d}n(1+\beta(\mathbb{R}^{d})+\varepsilon(n^{-1/d})).$
@@ 3.3 Local to global allocation
The following proposition formalizes and generalizes one of the steps in the proof of [33, Theorem 1(iii)] and is very important to deduce properties of global matchings via properties of matchings in finite windows. Recall that for a measure $\mu$, $\tilde{\mu}_{n}=\frac{n}{N}\mu\mathbf{1}_{\Lambda_{n}}$ (with $\tilde{\mu}_{n}=n\delta_{0}$ for $N=0$) and $\tilde{\mathcal{L}}_{n}:=\frac{1}{(2\pi)^{d}}\mathcal{L}\mathbf{1}_{\Lambda_{n}}$ are the renormalised samples with mass $n$ on $\Lambda_{n}.$
@@
Let $\mu,\nu$ be independent simple point processes with integrable RPCM and $w:[0,\infty)\to[0,\infty)$ be a continuous increasing weight function with $w(0)=0$. Assume $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$. Then, there exists an invariant coupling $M$ of $\mu$ and $\nu$ that is a.s. a matching and whose typical distance $X$ satisfies $\mathbf{E}w(X)<\infty$.
While the existence of a coupling of $\mu,\nu$ proceeds by a compactness argument, the fact that the coupling is a matching is an infinite version of the Birkhoff-von Neumann theorem ([53, Exercise 40]) which guarantees one-to-one correspondance in the case where $\mu,\nu$ are finite simple point processes with same number of points. For the infinite case, we exploit a recent result of [25]. However, a more explicit (but tedious) construction appears in a previous version of this work on arXiv.
Before proceeding with this proof, we need to recall some results on vague convergence from [22, Section 11]. For a measure $m$ on $\mathbb{R}^{d}$ and a measurable function $f:\mathbb{R}^{d}\to\mathbb{R}$, we denote $m(f):=\int f\,\mathrm{d}m$ whenever the integral is well-defined. Given measures $m,m_{n},n\geqslant 1$ on $\mathbb{R}^{d}$, say that $m_{n}$ converges vaguely to $m$, denoted by $m_{n}\xrightarrow[]{v}m\;$, if $m_{n}(f)\to m(f)$ for all $f\in C_{c}(\mathbb{R}^{d})$, the space of compactly supported (bounded) continuous functions on $\mathbb{R}^{d}$. For random measures $M_{n},M$, say $M_{n}$ converges to $M$ vaguely in law if $M_{n}(f)\stackrel{{\scriptstyle(d)}}{{\to}}M(f)$ for compactly supported (bounded) continuous $f$, where $\stackrel{{\scriptstyle(d)}}{{\to}}$ denotes convergence in law. Recall from [22, Proposition 11.1.VI] that if a sequence of random stationary measures $(M_{n})_{n}$ satisfies
$\sup_{n}\mathbf{E}M_{n}(A)<\infty$ (44)
@@
The first two steps are actually valid for any stationary random non-negative measure with finite intensity while the third step exploits the assumption of simple point processes, we believe that this can be avoided.
First step: Existence of an invariant infinite coupling $M$:
Let $\epsilon>0$. By definition (10), for every realization there is a transport plan $m_{n}$ from $\tilde{\mu}_{n}$ to $\tilde{\nu}_{n}$ such that
$\int_{\Lambda_{n}^{2}}w(d_{n}(x,y))m_{n}(\mathrm{d}(x,y))\leq\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})+\epsilon n.$
Thus by assumption $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})\leq cn$, we have that
@@
$m_{n}(\cdot,\mathbb{R}^{d})=\tilde{\nu}_{n}$, $m_{n}(\mathbb{R}^{d},\cdot)=\tilde{\mu}_{n}$ where $d_{n}$ is the toroidal distance on $\Lambda_{n}$ and we have denoted the new constant as well by $c$, with an abuse of notation. Suppose we tile $\mathbb{R}^{d}$ with randomly shifted copies of $\Lambda_{n}$ and place $\tilde{\mu}_{n},\tilde{\nu}_{n}$ in each of them and match by $m_{n}$ in each copy. We call this transport map $M_{n}$. Formally,
$M_{n}(A,B)=\sum_{z\in\mathbb{Z}^{d}}m_{n}\big((A-U_{n}+{2\pi}n^{1/d}z),(B-U_{n}+{2\pi}n^{1/d}z)\big),$ (46)
for $A,B\subset\mathbb{R}^{d}$ and where $U_{n}$ is uniform in $\Lambda_{n}$ and independent of $\mu$ and $\nu$. Observe that the translation-invariance of $\sum_{z\in\mathbb{Z}^{d}}\delta_{{2\pi}n^{1/d}z-U_{n}}$ shows that $M_{n}$’s are invariant or equivariant, i.e.
${M_{n}(A+x,B+x)\stackrel{{\scriptstyle(d)}}{{=}}M_{n}(A,B)},$ (47)
for all $A,B\subset\mathbb{R}^{d}$ and $x\in\mathbb{R}^{d}$. Note that $M_{n}(A,B)=0$ if $A,B$ are contained in different shifted copies.
Note that $\mathbf{E}M_{n}(A,B)\leqslant\mathbf{E}M_{n}(A,\mathbb{R}^{d})$ and $\mathbf{E}M_{n}(A,\mathbb{R}^{d})$ is translation invariant in $A$. Thus $\mathbf{E}M_{n}(A,\mathbb{R}^{d})=c_{n}|A|$ for some $c_{n}\in[0,\infty]$. For $A=\Lambda_{n}$, we have that $\mathbf{E}M_{n}(A,\mathbb{R}^{d})=\mathbf{E}\tilde{\mu}_{n}(\Lambda_{n})=n$ and so we deduce that $c_{n}=1$. Similarly, we can deduce that $\mathbf{E}M_{n}(\mathbb{R}^{d},B)=|B|$. Hence, we have that $\mathbf{E}M_{n}(A,B)\leqslant\min(|A|,|B|)$.
In particular, $\mathbf{E}M_{n}(K)<\infty$ for bounded $K\subset\mathbb{R}^{d}\times\mathbb{R}^{d}$. Thus, appealing to (44), the sequence $M_{n}$ is tight in vague topology and has a subsequential distributional limit under vague topology, say $M_{n^{\prime}}\stackrel{{\scriptstyle(d)}}{{\to}}M$. For a compactly supported continuous function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{+}$ and $x\in\mathbb{R}^{d}$, define $f_{x}(\cdot,\cdot):=f(x+\cdot,x+\cdot)$, also a bounded continuous function. By invariance of $M_{n}$’s, we have that $M_{n}(f_{x})\stackrel{{\scriptstyle(d)}}{{=}}M_{n}(f)$ and hence the subsequential convergence argued above gives that $M_{n}(f_{x})\to M(f)$ and $M_{n}(f)\to M(f)$ as $n\to\infty$. Thus $M(f_{x})\stackrel{{\scriptstyle(d)}}{{=}}M(f)$, which gives that $M$ is invariant.
@@
Recalling (3) and (4), to show that the typical transport distance under $M$ has finite cost means that
$\mathbf{E}\int\int w(\|x-y\|)\mathbf{1}[x\in\Lambda_{1}]M(\mathrm{d}(x,y))<\infty.$ (48)
Denote $\tilde{d}_{n}(x,y)$ to be the distance which is 0 if $x,y$ are in different randomly shifted copies of $\Lambda_{n}$, and the toric distance of the copy otherwise i.e.,
$\tilde{d}_{n}(x,y):=\sum_{z\in\mathbb{Z}^{d}}d_{n}(x-U_{n}-{2\pi}n^{1/d}z,y-U_{n}-{2\pi}n^{1/d}z).$
So $\mathbf{E}[\int\int w(\tilde{d}_{n}(x,y))\mathbf{1}[x\in A]M_{n}(\mathrm{d}(x,y))]$ is translation invariant in $A$ and hence its intensity is either infinity or of the form $\textrm{const}\cdot\mathcal{L}$, and for $A=\Lambda_{n}$, by (45)
@@
Thus we obtain that, there exists a constant $c$ such that for all AA
@@
Now let $A=\Lambda_{1}$ and $B$ be a bounded set. Let $n_{0}$ be such that $\sup_{x\in\Lambda_{1},y\in B}\|x-y\|<n^{1/d}/2$ for all $n\geqslant n_{0}$. Hence for $x\in\Lambda_{1},y\in B$, either $x,y$ are in different shifted copies of $\Lambda_{n}$ and $M_{n}(\mathrm{d}x,\mathrm{d}y)=0$ (and $\tilde{d}_{n}(x,y)=0$), or they are in the same copy of $\Lambda_{n}$, but at distance $\frac{n^{1/d}}{2}$, hence $\tilde{d}_{n}(x,y)=\|x-y\|.$ In any case,
@@
$\mathbf{E}\int\int w(\|x-y\|)\mathbf{1}[x\in\Lambda_{1},y\in B]M_{n}(\mathrm{d}(x,y))$ $=\mathbf{E}\int\int w(\tilde{d}_{n}(x,y))\mathbf{1}[x\in\Lambda_{1},y\in B]M_{n}(\mathrm{d}(x,y))$
$\leqslant c\mathcal{L}(\Lambda_{1})<\infty.$
We have shown in the First Step that $M$ is the subsequential distributional limit under vague topology of $M_{n}$. For convenience, we now assume that $M_{n}\stackrel{{\scriptstyle(d)}}{{\to}}M$ in vague topology. For any continuous function $0\leqslant f(x,y)\leqslant\mathbf{1}[x\in\Lambda_{1},y\in B]$, the vague convergence in law yields that, with $\tilde{w}(x,y)=w(\|x-y\|)f(x,y)$,
$X_{n}:=\int\int\tilde{w}(x,y)\,M_{n}(\mathrm{d}(x,y))\stackrel{{\scriptstyle(d)}}{{\to}}X:=\int\int\tilde{w}(x,y)\,M(\mathrm{d}(x,y)).$
The upper bound then is a general feature of convergence in law, through a truncation argument for non-negative variables: for $T>0$ fixed, $\mathbf{E}X_{n}\wedge T\to\mathbf{E}X\wedge T$, and $\mathbf{E}X_{n}\wedge T\leqslant c\mathcal{L}(A)$, so that with monotone convergence, we derive
$\mathbf{E}\int\int\tilde{w}(x,y)\,M(\mathrm{d}(x,y))\leqslant c\mathcal{L}(\Lambda_{1})<\infty.$
Then letting $f(x,y)$ increase pointwise to $\mathbf{1}[x\in\Lambda_{1}]$ and using the monotone convergence theorem again gives that
$\mathbf{E}\int\int w(\|x-y\|)\mathbf{1}[x\in\Lambda_{1}]M(\mathrm{d}(x,y))\leqslant c\mathcal{L}(\Lambda_{1})<\infty$
as needed to verify finite typical cost in (48).
@@
We will now show that $M$ is a transport plan between independent copies of $\mu$ and $\nu$ defined on a suitable probability space.